OpenStudy (anonymous):

the product of three intergers is 90. the second number is twice the first number. the third number is two more then the first number. what are the three numbers

OpenStudy (anonymous):

x stands for 1st number y stands for 2nd number z stands for 3rd number x + y + z = 90 (the product of 3 integers is 90) y = 2x (the second number is twice the first number) z = x + 2 (the third number is two more then the first number) Now that you have gotten your equations, you can substitute 2x in for y and x + 2 in for z in the first equation. Try it....you got this.

OpenStudy (skullpatrol):

xyz = 90 (the PRODUCT of 3 integers is 90)

OpenStudy (anonymous):

skullpatrol is correct....the first equation is wrong...my bad...product is the result of a multiplication problem....so the first equation would be xyz = 90.